OpenStudy (anonymous):

Solve this 15x + 8y = $214

OpenStudy (anonymous):

To solve this, rewrite the equation in terms of y, that is, get y on a side by itself.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

So if we start with 15x + 8y = 214, let's isolate the 8y. To isolate it, let's subtract 15x from both sides. 15x + 8y = 214 -15x - 15x We get 8y = 214 - 15x

OpenStudy (anonymous):

Do you understand what I did? Whatever we do to one side, we must do to the other.

OpenStudy (anonymous):

x=? y=?

OpenStudy (anonymous):

In this case, all you would be able to find are maximums of x and y. You will not be able to evaluate completely.

OpenStudy (anonymous):

They are asking for a simplified equation.

OpenStudy (anonymous):

x=10 y=8

OpenStudy (anonymous):

That is only one set of possible solutions. There are an infinite amount of solutions.

OpenStudy (amistre64):

the solution is actually a line on a graph and not a single set of numbers.